OpenStudy (anonymous):

If your very smart, this is like simple math for you then. xD (Find the Value of X)

OpenStudy (anonymous):

Find the value of x. The diagram is not drawn to scale. I really dont know what to after my steps. In picture, the pentagon looks like so and all i did was 540 - the sum of the other angles. But after than im lost. Help?

OpenStudy (anonymous):

*you're

OpenStudy (anonymous):

see? thats why i need smart people

OpenStudy (anonymous):

first of all, may I ask you if you know what the summation of the internal angles of a pentagon is? well it's 180(n-2) n being the number of sides

OpenStudy (anonymous):

and 5-2 x 180 or 3 x 180 = 540. i think.

OpenStudy (anonymous):

so for a pentagon, there are 5 sides. substituting to the equation 180(5 - 2) = 180(3) = 540 so the summation of the angles inside your pentagon is 540.

OpenStudy (anonymous):

im with you so far.

OpenStudy (anonymous):

I just saw the lower part of your work and it's okay. Just continue your work.

OpenStudy (anonymous):

but how? i know that the missing angles together equal 190. but i dont know how to figure out X to make 190.

OpenStudy (anonymous):

just add the 2x and the 2x and remove the 10 from the side with the variables by ussing it's additive inverse on both sides 2x + 2x +(10 - 10) = 190 -10 4x = 180 x = 45

OpenStudy (anonymous):

Then, substitute afterwards

OpenStudy (anonymous):

is it clear?

OpenStudy (anonymous):

okay. i saw it now. Brain Dead. forgot about -10 on each side. so.... x = 45 and the angles are a 90 degree angle and a 100 degree angle. thanks man.

OpenStudy (anonymous):

No problem